A university invests $600,000 at simple interest, part at 6%, half that amount at 4.5%, and the rest at 3.5%. What is the most that the university can invest at 3.5% and be guaranteed $24,600 in interest per year?
Answer by lwsshak3(11628)

A university invests $600,000 at simple interest, part at 6%, half that amount at 4.5%, and the rest at 3.5%. What is the most that the university can invest at 3.5% and be guaranteed $24,600 in interest per year?
***
let x=amt invested at 4.5%
2x=amt invested at 6%
600,000-3x=amt invested at 3.5%
..
4.5%x+6%2x+3.5%(600,000-3x)=24,600
.045x+.12x+21000-.105x=24600
0.06x=3600
x=60000
600000-3x=420000
most that the university can invest at 3.5% = $420,000